Week of 060612

Open Actions from last week:
  • Add OPS VO to FTS (Gavin)
  • Intervention Monday morning - FTS, LFC, CASTOR, SFT, SAM, GRIDVIEW - DONE
Chair: James

Monday:

Log: R-GMA problems over the weekend

New Actions:

Discussion:

  • R-GMA problem was dsue to to a memory problem on registry - RAL operations team to fix.

Tuesday:

Log:

New Actions:

Discussion:

Wednesday

Log: Gridview stopped showing data from 10pm last night

Actions:

  • Investigate Gridview/R-GMA problem (Gridview/R-GMA teams) - DONE
  • check LFC_DB_ERROR on lfc005 and pass to developers (James) - DONE

Discussion:

  • GGUS tickets went through - this will be easier when CERN ROC is in Remedy.
  • FIO can take 20 nodes out of SC4 cluster when they need them ( we assumed they already had !)

Thursday

Log: LFC_ACTIVE_CONN on lfc001.

Actions:

Discussion:

  • security upgrades to come for SLC3/4

Friday

Log:

Actions:

Discussion:

Edit | Attach | Watch | Print version | History: r5 < r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r5 - 2006-06-20 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback